OpenStudy (anonymous):

Solve by factoring and using the principle of zero products. x2 - x = 12

OpenStudy (campbell_st):

x^2 - x - 12= 0 find the factors of -12 that add to -1 (x -4)(x+3)=0

OpenStudy (anonymous):

Factor by grouping. x3 + 5x2 - 2x - 10

OpenStudy (campbell_st):

x(x^2 -2) + 5(x^2 - 2) = (x+5)(x^2 -2) this is grouping in pairs

OpenStudy (anonymous):

thanx Factor by grouping. x3 - 4x2 - 8x + 32

OpenStudy (campbell_st):

x(x -8) -4(x^2 -8) = (x -4)(x^2 - 8)

OpenStudy (campbell_st):

easy... find 2 factors of -48 that add to 2.... good luck